OpenStudy (anonymous):

A 16 ounce can of mixed nuts costs $5.82, but peanuts cost only $.25 per ounce. The can contains 7 ounces of peanuts and 9 ounces of other nuts. What is the cost per ounce of the other nuts?

OpenStudy (dtan5457):

.25 times 7. 5.82-1.75. $4.07 left for the other nuts, divide that by 9.

OpenStudy (triciaal):

peanuts cost 0.25/oz and you have 7 ozs. so the cost is ? 7x + 9 y = 5.82 x = cost of peanuts = 0.25 and y = cost of other nuts y = (5.82 - 7*0.25) / 9 =